# Adaptive Canonicalization With Application to Invariant Anisotropic Geometric Networks

### v1

@@ 1 Introduction
Equivariant machine learning [Gerken et al., 2023, Villar et al., 2021, Han et al., 2022, Keriven and Peyré, 2019] has been accentuated in geometric representation learning [Bronstein et al., 2017], motivated by the need to build models that respect symmetry inherent in data. For example, permutation equivariance in graphs [Gilmer et al., 2017, Zaheer et al., 2017, Xu et al., 2018], translation equivariance in images [LeCun and Bengio, 1998, Cohen and Welling, 2016a], and SO(3) or SE(3) equivariance for 3D objects and molecules [Thomas et al., 2018, Fuchs et al., 2020, Batzner et al., 2022, Satorras et al., 2021]. The symmetry is built into the method so that transforming the input induces a predictable transformation of the output. This inductive bias reduces sample complexity, curbs overfitting to arbitrary poses, and often improves robustness on distribution shifts where the same object appears in a different orientation or ordering [Kondor and Trivedi, 2018, Wang et al., 2022, Park et al., 2022, Bronstein et al., 2021, Bietti and Mairal, 2019, Kaba and Ravanbakhsh, 2023].
There are three principal approaches to handling symmetry in machine learning. The first involves designing equivariant architectures [Cohen and Welling, 2016b, Weiler et al., 2018a, Weiler and Cesa, 2019, Geiger and Smidt, 2022, Maron et al., 2019, Lippmann et al., 2024]: neural network layers are constructed to commute with the symmetry. The second approach is data augmentation, where each datapoint is presented to the model at an arbitrary pose [Chen et al., 2020, Brandstetter et al., 2022]. The third strategy is canonicalization [Kaba et al., 2023, Ma et al., 2023, 2024, Lim et al., 2022, 2023, Mondal et al., 2023, Lawrence et al., 2025, Sareen et al., 2025, Luo et al., 2022]: each input is mapped to a standard form and then processed by a non-equivariant network. Another less common approach to equivariant machine learning is frame averaging [Puny et al., 2021], which averages the network’s output over a set of input transformations.
A well-known problem in canonicalization is that in many cases it unavoidably leads to an end-to-end architecture which is discontinuous with respect to the input [Dym et al., 2024, Zhang et al., 2019a, Lim et al., 2022]. This inevitably leads to problems in stability during training and in generalization, as very similar inputs can lead to very different outputs [Dym et al., 2024, Tahmasebi and Jegelka, 2025a, b]. Moreover, the discontinuity of the network makes universal approximation properties less natural, as one approximates continuous symmetry preserving functions with discontinuous networks [Dym et al., 2024, Kaba et al., 2023, Wagstaff et al., 2022]. See App. A for extended related work.
@@
To motivate this approach, consider a neural network $\Psi$ which, by virtue of being symmetriless, may perform better on some orientations of $x$ than others. For illustration, it is easier for humans to detect an image as a horse if its limbs point downwards. Suppose that $\Psi(x)=(\Psi_{j}(x))_{j=1}^{3}=(\Psi_{dog}(x),\Psi_{cat}(x),\Psi_{horse}(x))$ is a sequence of binary classifiers with values in $[0,1]$ each. The output of $\Psi$ is defined to be the class with highest probability. Suppose moreover that for each $j$, the network $\Psi_{j}$ is granted the ability to rotated $x$ freely, and probe the output $\Psi_{j}(\pi(\alpha)x)$ for each $\alpha$. The network then chooses the orientation $\alpha_{*}$ such that $\Psi_{j}(\pi(\alpha_{*})x)$ is maximized. As an analogy, one can imagine an image on a piece of paper being handed at a random orientation to a person with visual system $\Psi$. To detect if there is a horse in the image, the person would rotate the paper, searching for an orientation which looks like a horse. Namely, if there is an orientation $\alpha_{*}$ where $\Psi_{horse}(\pi(\alpha_{*})x)$ is high then there is a horse in the image, and otherwise there is non. This process would be repeated for all other classes, and eventually the image would be classified as the $j_{*}$ such that $\max_{\alpha}\Psi_{j_{*}}(\pi(\alpha)x)$ is greater than $\max_{\alpha}\Psi_{j}(\pi(\alpha)x)$ for all other $j\neq j_{*}$. This is the process that we call prior maximization adaptive canonicalization.
@@ 2.1 Basic Definitions and Background
The function that maps each input $x$ to the output $f(x)$ is denoted by $x\mapsto f(x)$. The free variable of a univariate function is denoted by $(\cdot)$, and by $(\cdot,\cdot\cdot)$ for a function of two variables. For example, the function $(x,y)\mapsto\sin(x)\exp(y)$ is also denoted by $\sin(\cdot)\exp(\cdot\cdot)$. We denote the infinity norm of $x=(x_{d})_{d=1}^{D}\in\mathbb{R}^{D}$ by $\left\lvert x\right\rvert:=\max_{1\leq d\leq D}\left\lvert x_{d}\right\rvert$. We define the infinity norm of a continuous function $f:\mathcal{K}\rightarrow\mathbb{R}^{D}$ over a topological space $\mathcal{K}$ by $\|f\|_{\infty}=\sup_{x\in\mathcal{K}}\left\lvert f(x)\right\rvert$. If $\|f-y\|_{\infty}<\epsilon$ we say that $y$ approximates $f$ uniformly up to error $\epsilon$. The set of all subsets of a set $\mathcal{K}$, i.e., the power set, is denoted by $2^{\mathcal{K}}$. When defining general metric spaces, we allow the distance between points to be $\infty$. This does not affect most of the common properties of metric spaces (see Burago et al. [2001]).
@@
  • A function $f:\mathcal{K}\rightarrow\mathbb{R}^{D}$ is said to vanish at infinity is for every $\epsilon>0$ there exists a compact set $K\subset\mathcal{K}$ such that $\left\lvert f(x)\right\rvert<\epsilon$ for every $x\in\mathcal{K}\setminus K$.
@@
In the above definition, we interpret $\mathcal{N}(\mathcal{K},\mathbb{R}^{D})$ as a space of neural networks. A UAT is hence any theorem which shows that some set of neural networks is a universal approximator. Two examples of UATs are: 1) multilayer perceptrons (MLP) are universal approximators of $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ for compact subset $\mathcal{K}$ of the Euclidean space ${\mathbb{R}}^{D}$ [Hornik et al., 1989, Cybenko, 1989], and 2) DeepSets [Zaheer et al., 2017] are universal approximators of continuous functions from multi-sets to ${\mathbb{R}}^{d}$. See App. B for more details.
@@ 2.3 Prior Maximization Adaptive Canonicalization
When used for classification, we interpret each output channel $f_{d}\circ\rho_{f_{d}}^{d}(g)\in[0,1]$ as a binary classifier, i.e., representing the probability of $g$ being in class $d$ vs. not being in class $d$. This multiclass classification setting is called one vs. rest, where a standard loss is a sum of $D$ binary cross-entropies [Rifkin and Klautau, 2004, Galar et al., 2011, Allwein et al., 2000].
@@ 3.1 Anisotropic nonlinear spectral filters
In anisotropic nonlinear spectral filters (A-NLSF), we consider a symmetryless neural network $\Psi$ that operates on the space of spectral coefficients of the signal, e.g., a multilayer perceptron (MLP). To define $\Psi$ consistently, over a Euclidean space of fixed dimension, we extend or truncate the sequence of spectral coefficients as follows. Let $J_{1},\ldots,J_{B}$ be a predefined sequence of integers, and denote $J=\sum_{k}J_{k}$. We define $P$ as the mapping that takes $({\mathbf{X}},\mathbf{S})$ as input, first compute the spectral coefficients $\big(C_{k}({\mathbf{X}},\mathbf{S})\big)_{k=1}^{B}$, and then truncates or pads with zeros each $C_{k}({\mathbf{X}},\mathbf{S})$ to be in ${\mathbb{R}}^{J_{k}\times T}$. Namely, $P({\mathbf{X}},\mathbf{S})=\big(P_{k}({\mathbf{X}},\mathbf{S})\in{\mathbb{R}}^{J_{k}\times T}\big)_{k=1}^{B}\in{\mathbb{R}}^{J\times T}$. Here, the matrix $\emptyset$ is padded to the zero matrix $\mathbf{0}\in{\mathbb{R}}^{J_{k}\times T}$. Hence, the symmetryless network $\Psi$ maps ${\mathbb{R}}^{J\times T}$ to ${\mathbb{R}}^{D}$. A more detailed derivation of the construction is given in App. E.1.
@@ 3.2 Anisotropic Point Cloud Networks
Here, we present a point cloud network which is a combination of an equivariant network with adaptive canonicalization. Namely, we consider a permutation invariant network $\Psi$ like DeepSet [Zaheer et al., 2017] or DGCNN [Wang et al., 2019], and to attain 3D rotation invariance in addition we incorporate prior maximization. Together, the method is invariant both to permutations and rotations. We call this method anisotropic since $\Psi$ does not respect the rotation symmetries, and is hence sensitive to directions in the $x,y,z$ space.
@@
In App. B.2 we show that the quotient space has a natural metric. We hence take $\mathcal{G}=\mathcal{K}$ consisting of a single galaxy $\mathcal{G}_{N}=({\mathbb{R}}^{N\times J}/\sim)$. We moreover show in App. B.2 that any universal approximator of permutation invariant functions in $C_{0}({\mathbb{R}}^{N\times J},{\mathbb{R}}^{D})$, e.g., DeepSet [Zaheer et al., 2017], canonically gives a universal approximator of general continuous functions in $C_{0}({\mathbb{R}}^{N\times J}/\sim,{\mathbb{R}}^{D})$.
The symmetry in our adaptive canonicalization is 3D rotations $\pi(u)\in{\mathbb{R}}^{3\times 3}$, where $u$ is in the rotation group $\mathcal{SO}(3)$. Namely, we consider $J=3$, and rotated the rows of ${\mathbf{X}}\in{\mathbb{R}}^{N\times 3}$ via ${\mathbf{X}}\pi(u^{-1})$. We take $P$ as the identity. Note that this construction can be easily extended to multisets of arbitrary sizes, by considering the galaxies $\mathcal{G}_{N}={\mathbb{R}}^{N\times 3}/\sim$ and groups $\mathcal{U}_{N}=\mathcal{S}_{N}$ for all $N\in\mathbb{N}$. For details on the theoretical construction see App. B.2, and for details on the architecture App. E.2.
@@ 4 Experiments
We evaluate the anisotropic nonlinear spectral filters (Sec. 3.1) on toy problems and graph classification, and test the anisotropic point cloud network (Sec. 3.2) on point cloud classification. The experimental details, including experimental setups and hyperparameters, are in App. F.
@@ 4.1 Experimental Evaluation of Anisotropic Nonlinear Spectral Filters
We compare A-NLSF with the following baselines: (i) MLP, (ii) GCN [Kipf, 2016], (iii) GAT [Veličković et al., 2017], (iv) GIN [Xu et al., 2018], (v) ChebNet [Defferrard et al., 2016a], (vi) NLSF [Lin et al., 2024a], and (vii) S2GNN [Geisler et al., 2024]. In addition, we test canonicalization baselines by combining GIN with frame averaging (FA) [Puny et al., 2021] and orthogonalized axis projection (OAP) [Ma et al., 2024]. Tab. LABEL:tab:medium_size_graph_classification reports the classification results. We see that competing methods remain at chance level while A-NLSF achieves high accuracy by adaptively resolving ambiguities, showing the advantage for orientation-sensitive learning on disjoint supports.
@@
We further evaluate A-NLSF on graph classification benchmarks from TUDataset [Morris et al., 2020]: MUTAG, PTC, ENZYMES, PROTEINS, and NCI1, and follow the experimental setup [Ma et al., 2019, Ying et al., 2018, Zhang et al., 2019b] (see App. F). We compare with the same baselines as in the grid signal orientation tasks. Tab. LABEL:tab:medium_size_graph_classification summarizes the classification performance. Canonicalization baselines generally improve over GIN. Notably, we observe that A-NLSF outperforms competing baselines, suggesting that our AC provides more informative representations compared to a fixed, precomputed canonical form or isotropic filters.
@@
To further assess the effectiveness of A-NLSF, we evaluate on large-scale molecular and protein benchmarks from Open Graph Benchmark (OGB) [Hu et al., 2020]: ogbg-molhiv, ogbg-molpcba, and ogbg-ppa. We compare with GCN, GIN, GatedGCN [Bresson and Laurent, 2017], PNA [Corso et al., 2020], GraphTrans [Wu et al., 2021], SAT [Chen et al., 2022], GPS [Rampášek et al., 2022], SAN [Kreuzer et al., 2021], and the canonicalization method OAP+GatedGCN. The molecular classification results are reported in Tab. 2, showing that our method constructs robust molecular and protein representations, leading to improved generalization in classification.

### v2

@@ 1 Introduction
Equivariant machine learning [gerken2023geometric, villar2021scalars, han2022geometrically, keriven2019universal] has been accentuated in geometric representation learning [bronstein2017geometric], motivated by the need to build models that respect symmetry inherent in data. For example, permutation equivariance in graphs [gilmer2017neural, zaheer2017deep, xu2018powerful], translation equivariance in images [lecun1998convolutional, cohen2016group], and SO(3) or SE(3) equivariance for 3D objects and molecules [thomas2018tensor, fuchs2020se, batzner20223, satorras2021n]. The symmetry is built into the method so that transforming the input induces a predictable transformation of the output. This inductive bias reduces sample complexity, curbs overfitting to arbitrary poses, and often improves robustness on distribution shifts where the same object appears in a different orientation or ordering [kondor2018generalization, wang2022approximately, park2022learning, bronstein2021geometric, bietti2019group, kaba2023symmetry].
There are three principal approaches to handling symmetry in machine learning. The first involves designing equivariant architectures [cohen2016steerable, weiler20183d, weiler2019general, geiger2022e3nn, maron2018invariant, lippmann2024beyond]: neural network layers are constructed to commute with the symmetry. The second approach is data augmentation, where each datapoint is presented to the model at an arbitrary pose [chen2020group, brandstetter2022lie]. The third strategy is canonicalization [kaba2023equivariance, ma2023laplacian, ma2024canonicalization, lim2022sign, lim2023expressive, mondal2023equivariant, lawrence2025improving, sareen2025symmetry, luo2022equivariant]: each input is mapped to a standard form and then processed by a non-equivariant network. Another common approach to equivariant machine learning is frame averaging [puny2021frame], which averages the network’s output over a set of input transformations.
A well-known problem in canonicalization is that in many cases it unavoidably leads to an end-to-end architecture which is discontinuous with respect to the input [dym2024equivariant, zhang2019fspool, lim2022sign]. This inevitably leads to problems in stability during training and in generalization, as very similar inputs can lead to very different outputs [dym2024equivariant, tahmasebi2025generalization, tahmasebi2025regularity]. Moreover, the discontinuity of the network makes universal approximation properties less natural, as one approximates continuous symmetry preserving functions with discontinuous networks [dym2024equivariant, kaba2023equivariance, wagstaff2022universal].
@@
To motivate this approach, consider a neural network $\Psi$ which, by virtue of being symmetriless, may perform better on some orientations of $x$ than others. For illustration, it is easier for humans to detect an image as a horse if its limbs point downwards. Suppose that $\Psi(x)=(\Psi_{j}(x))_{j=1}^{3}=(\Psi_{dog}(x),\Psi_{cat}(x),\Psi_{horse}(x))$ is a sequence of binary classifiers with values in $[0,1]$ each. The output of $\Psi$ is defined to be the class with highest probability. Suppose moreover that for each $j$, the network $\Psi_{j}$ is granted the ability to rotated $x$ freely, and probe the output $\Psi_{j}(\pi(\alpha)x)$ for each $\alpha$. The network then chooses the orientation $\alpha_{*}$ such that $\Psi_{j}(\pi(\alpha_{*})x)$ is maximized. As an analogy, one can imagine an image on a piece of paper being handed at a random orientation to a person with visual system $\Psi$. To detect if there is a horse in the image, the person would rotate the paper, searching for an orientation which looks like a horse. Namely, if there is an orientation $\alpha_{*}$ where $\Psi_{horse}(\pi(\alpha_{*})x)$ is high then there is a horse in the image, and otherwise there is non. This process would be repeated for all other classes, and eventually the image would be classified as the $j_{*}$ such that $\max_{\alpha}\Psi_{j_{*}}(\pi(\alpha)x)$ is greater than $\max_{\alpha}\Psi_{j}(\pi(\alpha)x)$ for all other $j\neq j_{*}$. This is the process that we call prior maximization adaptive canonicalization. This process is inspired by ideas from cognitive psychology, where the human visual system is believed to learn canonical mental models of objects and to discard redundant variation due to symmetries by mentally “rotating” perceived stimuli into alignment with these canonical views [shepard1971mental, cooper1973chronometric, tarr1989mental]. Our example of a person rotating a sheet of paper to recognize whether it contains a horse is directly inspired by this line of work: prior maximization adaptive canonicalization can be viewed as a neural analogue of mental rotation, where the network searches over transformations to align inputs with its learned canonical views. We note that some previous models in machine learning were also inspired by this process [palmer1981cannonical, harris2001object, graf2006coordinate, gomez2008memory, konkle2011canonical, risko2016cognitive, tacchetti2018invariant, schmidt2024tilt].
@@
2 Related Work
Canonicalization has been studied in several forms. For example, in computer vision and geometric deep learning methods, inputs are often first transformed into a standardized pose or reference frame before classification [lowe2004distinctive, jaderberg2015spatial]. More recent work formalizes this as an explicit canonicalization map feeding a downstream network [lim2022sign, ma2024canonicalization] or as energy-based canonicalization [kaba2023equivariance] in which one learns an energy over group elements and takes the minimizer as the canonical transformation. The latter has been further developed on symmetries defined by general Lie group actions [shumaylov2024lie]. Canonicalization has also been used for data alignment [mondal2023equivariant, schmidt2025robust] and for test-time optimization over transformations, where one searches over group actions to select a canonical representation before downstream inference [singhal2025test, schmidt2024tilt]. A related line of work is frame averaging [puny2020global], which averages a network’s output over a set of group transformations, and its extension to weighted frame averaging [dym2024equivariant], where each datapoint is equipped with a probability distribution over the group and averaging is performed with respect to this measure, yielding continuity guarantees. In our work, we instead study canonicalization as a function of both the input and the network, and we establish continuity guarantees for symmetry-preserving continuous functions realized by our construction. Moreover, our approach is not restricted to symmetries defined via group actions, and allows working with more general augmentations for transforming datapoints. We refer to App. A for further discussion and comparison with related work.
@@ 3.1 Basic Definitions and Background
The function that maps each input $x$ to the output $f(x)$ is denoted by $x\mapsto f(x)$. The free variable of a univariate function is denoted by $(\cdot)$, and by $(\cdot,\cdot\cdot)$ for a function of two variables. For example, the function $(x,y)\mapsto\sin(x)\exp(y)$ is also denoted by $\sin(\cdot)\exp(\cdot\cdot)$. We denote the infinity norm of $x=(x_{d})_{d=1}^{D}\in\mathbb{R}^{D}$ by $\left\lvert x\right\rvert:=\max_{1\leq d\leq D}\left\lvert x_{d}\right\rvert$. We define the infinity norm of a continuous function $f:\mathcal{K}\rightarrow\mathbb{R}^{D}$ over a topological space $\mathcal{K}$ by $\|f\|_{\infty}=\sup_{x\in\mathcal{K}}\left\lvert f(x)\right\rvert$. If $\|f-y\|_{\infty}<\epsilon$ we say that $y$ approximates $f$ uniformly up to error $\epsilon$. The set of all subsets of a set $\mathcal{K}$, i.e., the power set, is denoted by $2^{\mathcal{K}}$. When defining general metric spaces, we allow the distance between points to be $\infty$. This does not affect most of the common properties of metric spaces (see metric_book).
@@
  • A function $f:\mathcal{K}\rightarrow\mathbb{R}^{D}$ is said to vanish at infinity if for every $\epsilon>0$ there exists a compact set $K\subset\mathcal{K}$ such that $\left\lvert f(x)\right\rvert<\epsilon$ for every $x\in\mathcal{K}\setminus K$.
@@
In the above definition, we interpret $\mathcal{N}(\mathcal{K},\mathbb{R}^{D})$ as a space of neural networks. A UAT is hence any theorem which shows that some set of neural networks is a universal approximator. Two examples of UATs are: 1) multilayer perceptrons (MLP) are universal approximators of $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ for compact subset $\mathcal{K}$ of the Euclidean space ${\mathbb{R}}^{D}$ [hornik1989multilayer, cybenko1989approximation], and 2) DeepSets [zaheer2017deep] are universal approximators of continuous functions from multi-sets to ${\mathbb{R}}^{d}$. See App. B for more details.
@@ 3.3 Prior Maximization Adaptive Canonicalization
When used for classification, we interpret each output channel $f_{d}\circ\rho_{f_{d}}^{d}(g)\in[0,1]$ as a binary classifier, i.e., representing the probability of $g$ being in class $d$ vs. not being in class $d$. This multiclass classification setting is called one vs. rest, where a standard loss is a sum of $D$ binary cross-entropies [rifkin2004defense, galar2011overview, allwein2000reducing].
@@ 4.1 Anisotropic nonlinear spectral filters
In anisotropic nonlinear spectral filters (A-NLSF), we consider a symmetryless neural network $\Psi$ that operates on the space of spectral coefficients of the signal, e.g., a multilayer perceptron (MLP). To define $\Psi$ consistently, over a Euclidean space of fixed dimension, we extend or truncate the sequence of spectral coefficients as follows. Let $J_{1},\ldots,J_{B}$ be a predefined sequence of integers, and denote $J=\sum_{k}J_{k}$. We define $P$ as the mapping that takes $({\mathbf{X}},\mathbf{S})$ as input, first compute the spectral coefficients $\big(C_{k}({\mathbf{X}},\mathbf{S})\big)_{k=1}^{B}$, and then truncates or pads with zeros each $C_{k}({\mathbf{X}},\mathbf{S})$ to be in ${\mathbb{R}}^{J_{k}\times T}$. Namely, $P({\mathbf{X}},\mathbf{S})=\big(P_{k}({\mathbf{X}},\mathbf{S})\in{\mathbb{R}}^{J_{k}\times T}\big)_{k=1}^{B}\in{\mathbb{R}}^{J\times T}$. Here, the matrix $\emptyset$ is padded to the zero matrix $\mathbf{0}\in{\mathbb{R}}^{J_{k}\times T}$. Hence, the symmetryless network $\Psi$ maps ${\mathbb{R}}^{J\times T}$ to ${\mathbb{R}}^{D}$. A more detailed derivation of the construction is given in App. E.3.
@@ 4.2 Anisotropic Point Cloud Networks
Here, we present a point cloud network which is a combination of an equivariant network with adaptive canonicalization. Namely, we consider a permutation invariant network $\Psi$ like DeepSet [zaheer2017deep] or DGCNN [wang2019dynamic], and to attain 3D rotation invariance in addition we incorporate prior maximization. Together, the method is invariant both to permutations and rotations. We call this method anisotropic since $\Psi$ does not respect the rotation symmetries, and is hence sensitive to directions in the $x,y,z$ space.
@@
In App. B.2 we show that the quotient space has a natural metric. We hence take $\mathcal{G}=\mathcal{K}$ consisting of a single galaxy $\mathcal{G}_{N}=({\mathbb{R}}^{N\times J}/\sim)$. We moreover show in App. B.2 that any universal approximator of permutation invariant functions in $C_{0}({\mathbb{R}}^{N\times J},{\mathbb{R}}^{D})$, e.g., DeepSet [zaheer2017deep], canonically gives a universal approximator of general continuous functions in $C_{0}({\mathbb{R}}^{N\times J}/\sim,{\mathbb{R}}^{D})$.
The symmetry in our adaptive canonicalization is 3D rotations $\pi(u)\in{\mathbb{R}}^{3\times 3}$, where $u$ is in the rotation group $\mathcal{SO}(3)$. Namely, we consider $J=3$, and rotated the rows of ${\mathbf{X}}\in{\mathbb{R}}^{N\times 3}$ via ${\mathbf{X}}\pi(u^{-1})$. We take $P$ as the identity. Note that this construction can be easily extended to multisets of arbitrary sizes, by considering the galaxies $\mathcal{G}_{N}={\mathbb{R}}^{N\times 3}/\sim$ and groups $\mathcal{U}_{N}=\mathcal{S}_{N}$ for all $N\in\mathbb{N}$. For details on the theoretical construction see App. B.2, and for details on the architecture see App. E.4.
4.3 Additional Applications of Adaptive Canonicaluzation
Our adaptive canonicalization is a general framework and is not limited to the two applications considered above. For example, it can be instantiated for image truncation with a pretrained network, where the symmetry corresponds to different crops and prior maximization selects the crop on which the model is most confident (see App. E.5). Note that our setting and theoretical results also apply to pretrained networks on downstream tasks. Our formulation also accommodates several instances of continuous prior maximization, e.g., unbounded point clouds under rotations and continuous-to-discrete image settings with rotations and other image transformations (see App. C). Finally, to broadening the applicability of our approach beyond classification, we explore using the adaptive canonicalization mechanism for point cloud segmentation, further (see App. G.10).
@@ 5 Experiments
We evaluate the anisotropic nonlinear spectral filters (Sec. 4.1) on toy problems and graph classification, and test the anisotropic point cloud network (Sec. 4.2) on point cloud classification. The experimental details, including experimental setups and hyperparameters, are in App. F. Additional experiments, e.g., ablation study, are in App. G.
Maximization method.
We approximate the prior maximization by sampling a finite set of transformations from a probability measure on the transformation space, evaluating the prior for all candidates, and retaining the transformation that yields the largest prior value with the one-vs-rest classification objective. We then refine the selected candidate locally by running a few steps of gradient descent from the best sampled transformation (see App. E.2).
@@ 5.1 Experimental Evaluation of Anisotropic Nonlinear Spectral Filters
We compare A-NLSF with the following baselines: (i) MLP, (ii) GCN [kipf2016semi], (iii) GAT [velivckovic2017graph], (iv) GIN [xu2018powerful], (v) ChebNet [defferrard2016convolutional], (vi) NLSF [lin2024equivariant], and (vii) S2GNN [geisler2024spatio]. In addition, we test canonicalization baselines by combining GIN with frame averaging (FA) [puny2021frame] and orthogonalized axis projection (OAP) [ma2024canonicalization]. Tab. LABEL:tab:medium_size_graph_classification reports the classification results. We see that competing methods remain at chance level while A-NLSF achieves high accuracy by adaptively resolving ambiguities, showing the advantage for orientation-sensitive learning on disjoint supports.
@@
We further evaluate A-NLSF on graph classification benchmarks from TUDataset [morris2020tudataset]: MUTAG, PTC, ENZYMES, PROTEINS, and NCI1, and follow the experimental setup [ma2019graph, ying2018hierarchical, zhang2019hierarchical] (see App. F). We compare with the same baselines as in the grid signal orientation tasks. Tab. LABEL:tab:medium_size_graph_classification summarizes the classification performance. Canonicalization baselines generally improve over GIN. Notably, we observe that A-NLSF outperforms competing baselines, suggesting that our AC provides more informative representations compared to a fixed, precomputed canonical form or isotropic filters.
@@
To further assess the effectiveness of A-NLSF, we evaluate on large-scale molecular and protein benchmarks from Open Graph Benchmark (OGB) [hu2020open]: ogbg-molhiv, ogbg-molpcba, and ogbg-ppa. We compare with GCN, GIN, GatedGCN [bresson2017residual], PNA [corso2020principal], GraphTrans [wu2021representing], SAT [chen2022structure], GPS [rampavsek2022recipe], SAN [kreuzer2021rethinking], and the canonicalization method OAP+GatedGCN. The molecular classification results are reported in Tab. 2. We see that our method achieves consistent improvements across these datasets, leading to improved generalization in classification.